# Green Federated Learning: A new era of Green Aware AI

### v1

@@ 7.2. Smart Cities
(He et al., 2023) Provide a FL-based UAV clustering plan X Improve clustering stability and shorten service response times
(Nasr-Azadani et al., 2023) Make a recommendation to support ground wireless networks with FL-based heterogeneous UAVs X Increase response time, network throughput, and learning speed
(Wang et al., 2023a) Provide a FL-based framework for UAVs to handle resource constraints and heterogeneous data X Improve learning outcomes, reduce computing expenses, and raise UAV trajectory
(Liu et al., 2023a) By utilizing split learning and FL in a distributed fashion, suggest FL-based UAVs X Boost accuracy rate, convergence, and learning speed
(Yang et al., 2023) Provide an FL-based computing system that UAVs can use to track urban transit. X X Shorten response times for services, improve network performance, and learning speed.
(Hou et al., 2023) Make a security-conscious FL-based learning plan available for UAVs X Boost security while ensuring time and energy savings
(Fu et al., 2023a) Provide the FL convergence analysis without making any convexity assumptions and show how device scheduling affects the global gradients for UAV application X In comparison to current benchmarks, the suggested design considerably improves the tradeoff between completion time and test accuracy.
7.3. Smart Healthcare
As a new distributed collaborative AI paradigm, FL mainly appeals to smart healthcare since it allows several customers, like hospitals, to coordinate AI training without exchanging raw data. The latest FL ideas for smart healthcare encompass resource-conscious FL, secure and privacy-conscious FL, incentive-focused FL, and personalized FL (Nguyen et al., 2022). Table 9 discusses several emerging applications of FL in critical healthcare areas, such as COVID-19 detection, medical imaging, remote health monitoring, human activity recognition, and health data management.
@@ 7.3. Smart Healthcare
(Kang et al., 2021) Provide an FL and Blockchain-based work distribution plan for smart transportation systems X X The suggested methods have the potential to significantly increase federated edge learning security
(Sepasgozar and Pierre, 2022) Provide an FL method for Network Traffic Prediction (Fed-NTP) that trains the model locally using the Long-Short-Term Memory (LSTM) algorithm to accurately predict network traffic flow while maintaining privacy X Boost network traffic flow accuracy forecast rate, privacy, and security
(Liu et al., 2023b) Create a sophisticated multi-tier FL-based vehicle traffic forecasting system for intelligent transportation networks X Boost the prediction accuracy rate and system efficiency
(Xu et al., 2023) provides an asynchronous federated learning technique with a dynamic scaling factor that is based on blockchain Verify superior dependability, efficiency, and learning performance
(Yuan et al., 2023) Provide a FL-IoT solution for smart city traffic flow prediction X Increase forecast accuracy while cutting down on time expenditure
(Hijji et al., 2023) provide an intelligent hierarchical framework for maintaining road infrastructure Support large-scale real-time smart transportation systems with adaptive resource allocation
7.4. Unmanned Aerial Vehicles (UAVs)
@@
7.5. Smart Transportation
The Internet of Things (IoT), sensor technologies, and communication technologies have developed rapidly, spurring the growth of smart transportation systems (STS). However, because vehicle networks are dynamic, it is difficult to decide on vehicle behaviors in a timely and correct manner. Furthermore, there is always a risk to the security and privacy of vehicle information when mobile wireless connections are present. One possible method to improve the effectiveness and safety of intelligent transportation systems is federated learning (Zhang et al., 2023a). Although FL has been examined in great detail, more is needed to understand the communication and networking issues that arise when FL is used in densely populated but dynamic vehicle networks. The limited storage and communication capacities of individual vehicles hinder the timely training of an FL model in distributed vehicular networks (Sangdeh et al., 2023). Here, we examine a few common FL-based ITS systems, as shown in Table 11.
8. Discussions & Future Scope
@@
  • A way to effectively address the challenges described in Section 3 to FL frameworks for efficiency and effectiveness should be determined. By offering a fundamental foundation for efficient model learning, FL enables users to work together to train a global model with their own datasets. However, there is no guarantee of effectiveness due to statistical and system heterogeneity of data and client devices. To address statistical heterogeneity, data moderation, personalization, and cluster mechanisms have been widely developed in a variety of FL frameworks (Section 3.1). Nevertheless, the proposed techniques used in the FL frameworks frequently reduce accuracy or effectiveness in addressing statistical and system heterogeneity. Therefore, while addressing heterogeneity in the FL, it is essential to weigh the trade-off between accuracy and energy efficiency.
@@
  • Numerous methods have been put forth to increase federated learning’s communication effectiveness described in Section5. These methods mostly focus on reducing the size of the model updates that are sent or the frequency at which these updates are communicated. Nevertheless, the advantages of using such solutions frequently come at the expense of a reduction in the accuracy of the classification or forecast. To create effective and efficient communication models for federated learning, a thorough examination of the trade-off between accuracy preservation and communication efficiency should be done. More precisely, we require novel communication strategies that can demonstrate efficiency at the Pareto frontier—that is, strategies that can provide an accuracy level greater than any other strategy at the same communication budget and over a wide range of communication/accuracy profiles.
@@
  • Key issues that impact FL performance include communication and computation efficiency described in Section 3. To address these issues, it is required to reduce the communication rounds, use an efficient client selection algorithm, or compress the deep learning model. Several techniques to reduce the communication rounds, client selection, and model compression are proposed in the literature described in Section 5. However, there is a lack of evidence to maintain the tradeoff between accuracy, communication efficiency, and energy efficiency in a heterogeneous environment. Therefore, while addressing effectiveness in the FL, it is essential to weigh the trade-off between accuracy, communication efficiency, and energy efficiency.
@@
  • Energy and carbon footprint analysis is a key ingredient to propose an FL-based IoT application for a sustainable environment described in Section 6. In the literature, there is no evidence available for energy and carbon footprint analysis for FL-based IoT applications. Therefore, while maintaining effectiveness and efficiency, it is essential to analyze the energy and carbon footprint to validate green-aware AI. Hence, there is a need for numerical assessment of many communication- and computationally-efficient FL techniques for IoT applications in terms of energy use and carbon equivalent emissions.
9. Conclusion

### v2

@@ Introduction
[He et al.(2023)] Provide a FL-based UAV clustering plan X Improve clustering stability and shorten service response times
[Nasr-Azadani et al.(2023)] Make a recommendation to support ground wireless networks with FL-based heterogeneous UAVs X Increase response time, network throughput, and learning speed
[Wang et al.(2023a)] Provide a FL-based framework for UAVs to handle resource constraints and heterogeneous data X Improve learning outcomes, reduce computing expenses, and raise UAV trajectory
[Liu et al.(2023a)] By utilizing split learning and FL in a distributed fashion, suggest FL-based UAVs X Boost accuracy rate, convergence, and learning speed
[Yang et al.(2023)] Provide an FL-based computing system that UAVs can use to track urban transit. X X Shorten response times for services, improve network performance, and learning speed.
[Hou et al.(2023)] Make a security-conscious FL-based learning plan available for UAVs X Boost security while ensuring time and energy savings
[Fu et al.(2023a)] Provide the FL convergence analysis without making any convexity assumptions and show how device scheduling affects the global gradients for UAV application X In comparison to current benchmarks, the suggested design considerably improves the tradeoff between completion time and test accuracy.
Smart Healthcare
As a new distributed collaborative AI paradigm, FL mainly appeals to smart healthcare since it allows several customers, like hospitals, to coordinate AI training without exchanging raw data. The latest FL ideas for smart healthcare encompass resource-conscious FL, secure and privacy-conscious FL, incentive-focused FL, and personalized FL [Nguyen et al.(2022)]. Table 9 discusses several emerging applications of FL in critical healthcare areas, such as COVID-19 detection, medical imaging, remote health monitoring, human activity recognition, and health data management.
@@
[Kang et al.(2021)] Provide an FL and Blockchain-based work distribution plan for smart transportation systems X X The suggested methods have the potential to significantly increase federated edge learning security
[Sepasgozar and Pierre(2022)] Provide an FL method for Network Traffic Prediction (Fed-NTP) that trains the model locally using the Long-Short-Term Memory (LSTM) algorithm to accurately predict network traffic flow while maintaining privacy X Boost network traffic flow accuracy forecast rate, privacy, and security
[Liu et al.(2023b)] Create a sophisticated multi-tier FL-based vehicle traffic forecasting system for intelligent transportation networks X Boost the prediction accuracy rate and system efficiency
[Xu et al.(2023)] provides an asynchronous federated learning technique with a dynamic scaling factor that is based on blockchain Verify superior dependability, efficiency, and learning performance
[Yuan et al.(2023)] Provide a FL-IoT solution for smart city traffic flow prediction X Increase forecast accuracy while cutting down on time expenditure
[Hijji et al.(2023)] provide an intelligent hierarchical framework for maintaining road infrastructure Support large-scale real-time smart transportation systems with adaptive resource allocation
Unmanned Aerial Vehicles (UAVs)
@@
Smart Transportation
The Internet of Things (IoT), sensor technologies, and communication technologies have developed rapidly, spurring the growth of smart transportation systems (STS). However, because vehicle networks are dynamic, it is difficult to decide on vehicle behaviors in a timely and correct manner. Furthermore, there is always a risk to the security and privacy of vehicle information when mobile wireless connections are present. One possible method to improve the effectiveness and safety of intelligent transportation systems is federated learning [Zhang et al.(2023a)]. Although FL has been examined in great detail, more is needed to understand the communication and networking issues that arise when FL is used in densely populated but dynamic vehicle networks. The limited storage and communication capacities of individual vehicles hinder the timely training of an FL model in distributed vehicular networks [Sangdeh et al.(2023)]. Here, we examine a few common FL-based ITS systems, as shown in Table 11.
Discussions & Future Scope
@@
  • A way to effectively address the challenges described in Section Challenges and solutions in Federated Learning to FL frameworks for efficiency and effectiveness should be determined. By offering a fundamental foundation for efficient model learning, FL enables users to work together to train a global model with their own datasets. However, there is no guarantee of effectiveness due to statistical and system heterogeneity of data and client devices. To address statistical heterogeneity, data moderation, personalization, and cluster mechanisms have been widely developed in a variety of FL frameworks (Section Optimization of Effectiveness). Nevertheless, the proposed techniques used in the FL frameworks frequently reduce accuracy or effectiveness in addressing statistical and system heterogeneity. Therefore, while addressing heterogeneity in the FL, it is essential to weigh the trade-off between accuracy and energy efficiency.
@@
  • Numerous methods have been put forth to increase federated learning’s communication effectiveness described in SectionVariants of Federated Learning Frameworks. These methods mostly focus on reducing the size of the model updates that are sent or the frequency at which these updates are communicated. Nevertheless, the advantages of using such solutions frequently come at the expense of a reduction in the accuracy of the classification or forecast. To create effective and efficient communication models for federated learning, a thorough examination of the trade-off between accuracy preservation and communication efficiency should be done. More precisely, we require novel communication strategies that can demonstrate efficiency at the Pareto frontier—that is, strategies that can provide an accuracy level greater than any other strategy at the same communication budget and over a wide range of communication/accuracy profiles.
@@
  • Key issues that impact FL performance include communication and computation efficiency described in Section Challenges and solutions in Federated Learning. To address these issues, it is required to reduce the communication rounds, use an efficient client selection algorithm, or compress the deep learning model. Several techniques to reduce the communication rounds, client selection, and model compression are proposed in the literature described in Section Variants of Federated Learning Frameworks. However, there is a lack of evidence to maintain the tradeoff between accuracy, communication efficiency, and energy efficiency in a heterogeneous environment. Therefore, while addressing effectiveness in the FL, it is essential to weigh the trade-off between accuracy, communication efficiency, and energy efficiency.
@@
  • Energy and carbon footprint analysis is a key ingredient to propose an FL-based IoT application for a sustainable environment described in Section Energy and Carbon Footprint Study in Federated Learning. In the literature, there is no evidence available for energy and carbon footprint analysis for FL-based IoT applications. Therefore, while maintaining effectiveness and efficiency, it is essential to analyze the energy and carbon footprint to validate green-aware AI. Hence, there is a need for numerical assessment of many communication- and computationally-efficient FL techniques for IoT applications in terms of energy use and carbon equivalent emissions.
Conclusion